### v1

@@ Abstract
We study dynamic decentralized two-sided matching in which players may encounter unanticipated experiences. As they become aware of these experiences, they may change their preferences over players on the other side of the market. Consequently, they may get “divorced” and rematch again with other agents, which may lead to further unanticipated experiences etc. A matching is stable if there is absence of pairwise common belief in blocking. Stable matchings can be destabilized by unanticipated experiences. Yet, we show that there exist self-confirming outcomes that are stable and do not lead to further unanticipated experiences. We introduce a natural decentralized matching process that, at each period assigns probability $1-\varepsilon$ to the satisfaction of a mutual optimal blocking pair (if it exists) and picks any optimal blocking pair otherwise. The parameter $\varepsilon$ is interpreted as a friction of the matching market. We show that for any decentralized matching process, frictions are necessary for convergence to stability even without unawareness. Our process converges to self-confirming stable outcomes. Further, we allow for bilateral communication/flirting that changes the awareness and say that a matching is flirt-proof stable if there is absence of communication leading to pairwise common belief in blocking. We show that our natural decentralized matching process converges to flirt-proof self-confirming outcomes.
@@ 1 Introduction
Many matching problems such as whom to marry, which school to choose, which profession to enter, where to take up a residency etc. involve transformative experiences (Paul, 2014) that change who we are, our beliefs, the things and issues we care about, and our preferences. These preference changes have implications for the stability of matchings. For instance, in the US marriage market, 43% of ever-married couples were divorced or widowed, and 23% of married couples are remarried couples (Livingstone, 2014). During marriage, spouses may become aware of intolerable attributes of the partner that they were previously oblivious to it. Or they may experience events in the family and workplace such as addiction, unemployment, domestic violence, mental health issues etc., some of which can only be fully grasped once experienced. For instance, marriage often involves parenting, itself a transformative experience, with profound changes in preferences as spouses become parents. Svar and Verner (2008) studied relationships in Denmark and found a negative causal impact of children on relationship duration.
@@
In Section 7 we study welfare in our setting of unanticipated experiences, changing awareness, and endogenous preferences. Since we consider unawareness but perfect information, we work with ordinal preferences, which makes Pareto efficiency a natural welfare criterion just like in the case of matching under complete information. However, Pareto efficiency is not straightforward in settings with endogenous preferences such as ours. Even if there are no Pareto improvements for a matching before the matching is experienced, there might be Pareto improvements to the matching after it is experienced (and vice versa). We define a notion of persistent Pareto efficiency. We argue that self-confirming outcomes may not be persistent Pareto efficient but every persistent Pareto efficient outcome is a self-confirming outcome.
In the final Section 8, we discuss the effect of infidelity. We also discuss how players can be confused about other players in self-confirming outcomes leading to (stable) awareness of unawareness (Schipper, 2024). Finally, we discuss the related literature and point to potential extensions of our work.
@@ 2.3 Random Paths to Stability
Prior section showed that we need to go beyond deterministic processes in order to guarantee stable outcomes in decentralized matching markets. Roth and Vande Vate (1990) showed that we can reach stable outcomes when a process that randomly picks among blocking pairs. In particular, they showed that starting from an arbitrary matching $\mu$, there exist a finite sequence of matchings $\mu_{1},...,\mu_{k}$ such that $\mu=\mu_{1}$ and $\mu_{k}$ is stable, and for $\mu_{i}$, $i=1,...,k-1$, there is a blocking pair $(m_{i},w_{i})$ that if satisfied yields matching $\mu_{i+1}$. As a corollary, they show that the process of satisfying randomly chosen blocking pairs will converge to a stable matching with probability one.
@@
This observation has a profound economic corollary: Frictions that prevent the satisfaction of mutually optimal blocking pairs are necessary for any decentralized matching processes to reach stability. This observation runs counter to standard economic wisdom that frictions hamper the functioning of decentralized markets. In contrast, for decentralized matching markets, small frictions are necessary for reaching stable matchings. In Appendix A, we present another application of our cyclic example to entry in matching markets.
@@ 2.4 Arbitrarily Small Frictions are Enough
For the unperturbed process, both stable matchings and cycles formed by satisfying unique mutually optimal blocking pairs are absorbing sets. However, we argue that for an arbitrarily small $\varepsilon>0$, which captures the frictions that prevent the matching of mutually optimal blocking pairs, the perturbed process converges to a stable matching in finite time with probability 1. Given a marriage market $\langle M,W,(\succ_{i})_{i\in M\cup W}\rangle$, cycle(s) formed by satisfying unique mutually optimal blocking pairs at each step may or may not exist. If there is no such cycle, our claim is trivially true. If a cycle by satisfying the unique mutually optimal blocking pair at each step exists, conditional on entering such a cycle, the probability that the process stays in this cycle for exactly $k$ period is $(1-\varepsilon)^{k-1}\varepsilon$. Hence the probability that the process stays in this cycle forever is $\lim_{k\rightarrow\infty}(1-\varepsilon)^{k-1}\varepsilon=0$, which means that the process leaves the cycle in finite time with probability 1. After leaving the cycle, the process might come back to this cycle or enter another cycle. However, the same argument applies, and the process leaves the cycle again in finite time with probability 1. As soon as the process catches one of Ackermann et al. (2008)’s optimal blocking pairs, it is trapped in a basin of attraction of a stable matching. Note that by the $\varepsilon$-events, any optimal blocking pair has a strict positive probability. Thus, we can make use of the result by Ackermann et al. (2008) according to which for any unstable matching, there exist a finite sequence of satisfying optimal blocking pairs that leads to a stable matching. The process may potentially spend long periods in the cycle(s), but only the stable matchings are absorbing
@@ 3 Matching under Unawareness
We illustrate an unawareness structure with point-beliefs for two players satisfying Assumptions 1 and 2 in Figure 3. There are two characteristics of players or experiences with players of which players may or may not be unaware. Thus, we have four spaces. The left space $S_{1}$ models the situation in which only the first characteristic/experience is expressible, while the right space $S_{2}$ models the situations when only the second characteristic/experience is expressible. On the upmost space $S_{3}$ both characteristics/experiences are expressible. In the lowest space $S_{0}$ none of the characteristics/experiences are expressible. States model combinations of awareness of both agents in a consistent way as described by the type mapping and the conditions imposed on the type mapping. The type mapping of one player (i.e., the man) is is indicated with blue solid lines while the one of the other player is indicated with pink dash lines (i.e., the woman). For instance, at state $\omega_{10}$ the woman is aware of both characteristics/experiences, because her type mapping maps $\omega_{10}$ to itself, while the man is only aware of the first characteristic/experience, because his type mapping maps $\omega_{10}$ to $\omega_{4}$ in $S_{1}$. At $\omega_{4}$ both players are aware of the first characteristic/experience. Thus, at $\omega_{10}$ the man point-believes that the woman is aware of the first characteristics/experiences, which is consistent with his awareness at $\omega_{10}$ since he can only envision the first characteristic/experience but not the second at $\omega_{10}$. More interesting is for instance state $\omega_{24}$. At that state, the man is aware of the first characteristic/experience only while the woman is only aware of the second characteristics/experiences only because their type mappings map to $\omega_{3}$ and $\omega_{6}$, respectively. At $\omega_{3}$, the man correctly believes that the woman is unaware of characteristic/experience 1 while at $\omega_{6}$ the woman correctly believes that the man is unaware of characteristic/experience 2. This is indicated by the type mappings that map the respective states to $\omega_{0}$. We omit the projections in order not to clutter the figure further, but they should be clear. For instance, states $\omega_{13},\omega_{14}$ and $\omega_{15}$ all project to $\omega_{7}$. Assumptions 1 and 2 are easy to verify for instance in spaces $S_{1}$ and $S_{2}$ (albeit less obvious in $S_{3}$). Note that any unawareness structure for correct point-beliefs of two players and two characteristics/experiences satisfying Assumptions 1 and 2 must look like Figure 3. With more players, which we need in non-trivial matching games, or more characteristics/experiences, they become more complex. However, in our examples we will often just focus on the states that are relevant for the feature we aim to illustrate with the example.
@@ 4 Self-Confirming Stable Outcomes
Stability itself is not sufficient for a satisfactory solution because $\mu$ being stable at $\omega$ does not rule out that some players discover something in the matching $\mu$ at $\omega$ that changes their preferences and consequently destabilizes the previously stable matching so that they want to get divorced. To obtain a satisfactory solution, we also need that beliefs are stable given the matching.
@@
We define the matching process by transition probabilities on the product space $\Omega\times\mathsf{M}$. This warrants some explanations. Naturally, the process would proceed as follows: Given a current state $\omega$ and awareness/preferences at $\omega$, the current match $\mu$ may feature some blocking pairs w.r.t. preferences at $\omega$. Satisfying some blocking pair (i.e., preferably some mutual optimal blocking pair) would lead to another match $\mu^{\prime}$. At this match $\mu^{\prime}$ and state $\omega$, there may be discoveries leading to another state $\tau(\omega,\mu^{\prime})=\omega^{\prime}$. At this state and corresponding awareness/preferences, there might be blocking pairs. Satisfying a blocking pair may lead to yet another match $\mu^{\prime\prime}$ While in this explanation, we let states transit after the transition of the matches, the above formal description of the process involves a simultaneous transition of states and matches. It can be thought of just describing the process at every “even” period. This is well-defined because states move deterministically according to a finite state machine given the prior state and the new match. The process on $\Omega\times\mathsf{M}$ is not deterministic though because blocking pairs are selected randomly with $(1-\varepsilon)$ assigned to mutual best blocking pairs if they exist.
@@ 5 Allowing for Flirting
$H_{m}(\omega,\mu,w):=\{\omega^{\prime}\in S_{t_{m}(\omega)}:\omega^{\prime}\in CK_{m,w}([m\succ_{w}\mu(w)]\cap[w\succ_{m}\mu(m)]),S_{t_{w}(\omega^{\prime})}\trianglerighteq S_{t_{w}(\omega)}\}$
@@ 6 Does Divorce Improve Welfare?
Example 8 Suppose there are five men $M=\{m_{1},m_{2},m_{3},m_{4},m_{5}\}$ and five women $W=\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Consider the following profile of strict preferences given by the rank order lists:
@@
and initial matching $\mu_{1}=\begin{pmatrix}m_{1}&m_{2}&m_{3}&m_{4}&m_{5}\\ w_{1}&w_{2}&w_{3}&w_{4}&w_{5}\end{pmatrix}$, which is stable given the above preference profile.
@@
Suppose $m_{1}$ is enlightened and changes his preference such that his new rank order list is $w_{2},w_{1},w_{4},w_{3},w_{5}$. For simplicity, suppose that this is the only change in the players’ preferences along the entire process and that all players have correct beliefs about every other agents’ preferences. With $m_{1}$’s new preference, he blocks with $w_{2}$. We show that if $m_{1}$ rematches to $w_{2}$, the process of satisfying mutually optimal blocking pairs may lead to any of the three stable matchings, depending on which mutually optimal blocking pair is satisfied at stages 7 or 8 of the process (see Figure 12). Furthermore, in these stable matchings, $m_{1}$ can be matched to $w_{2}$ (in $\mu_{10_{A}}$), $w_{1}$ (in $\mu_{11_{B}}$), or $w_{4}$ (in $\mu_{9}$), which means that he may be better off, same, or worse off.
@@
7 Pareto Efficiency
The previous section motivates the question about how to evaluate welfare in our setting. Under complete information, Pareto efficiency is used as welfare criterion for NTU matching markets. Focusing on unawareness but perfect information allows us to consider ordinal preferences just as in the case of complete information. However, the extension of Pareto efficiency to our setting is not straightforward because unanticipated experiences raise awareness and potentially change preferences. We need a notion of Pareto efficiency for endogenous preferences. Focusing for simplicity on matching without flirting, consider the following notion of Pareto improvement: Given an outcome $(\mu,\omega)$, the matching $\mu^{\prime}\neq\mu$ is a Pareto improvement if for all $i\in M\cup W$, $\mu^{\prime}(i)\neq\mu(i)$ implies $\mu^{\prime}(i)\ {\succ_{i}}(t_{i}(\omega))\ \mu(i)$. In order to make sense of the definition recall that throughout the paper we assumed for simplicity that players have strict preferences. Thus, when a matching $\mu^{\prime}$ makes some players better off without making anybody worse off as compared to matching $\mu$, matching $\mu^{\prime}$ must make all players better off (compared to $\mu$) for which the partner under $\mu^{\prime}$ differs from their partner under matching $\mu$. While the definition of Pareto improvement is a straightforward extension of Pareto improvement under complete information, it seems too weak in our context of endogenously changing preferences. In some sense, it is an ex ante notion of Pareto improvement before matching $\mu$ has been experienced by the players. We may want to require that there is also no Pareto improvement after $\mu$ has been experienced at $\omega$, because awareness may have changed and with it preferences. That is, given an outcome $(\mu,\omega)$, say that matching $\mu^{\prime}\neq\mu$ is a $\mu$-experienced Pareto improvement if for all $i\in M\cup W$, $\mu^{\prime}(i)\neq\mu(i)$ implies $\mu^{\prime}(i)\ {\succ_{i}}(t_{i}(\tau(\omega,\mu))\ \mu(i)$. It should be clear that $\mu^{\prime}$ might be a Pareto improvement given $(\mu,\omega)$ but not a $\mu$-experienced Pareto improvement because $\mu^{\prime}$ might be worse for some players than $\mu$ after having experienced $\mu$ and “developed a taste for it.” Conversely, $\mu^{\prime}$ might be a $\mu$-experienced Pareto improvement but not a Pareto improvement because after experiencing $\mu$ players discover much to hate about it and prefer now $\mu^{\prime}$. Thus, either notion does not imply the other. Requiring both notions may still not be enough: There could be a Pareto improvement after the matching $\mu$ has been experienced at $\omega$ and experienced again at $\tau(\omega,\mu)$, etc.
To define a notion of Pareto efficiency as persistent absence of Pareto improvements, let for all $\omega\in\Omega$ and $\mu\in\mathsf{M}$, $\tau^{0}(\omega,\mu):=\omega$, $\tau^{1}(\omega,\mu):=\tau(\omega,\mu)$, and recursively for any $n=1,2,...$, $\tau^{n}(\omega,\mu):=\tau(\tau^{n-1}(\omega,\mu),\mu)$.
Definition 7
We say that outcome $(\mu,\omega)$ is persistent Pareto efficient if for any $n=0,1,...$, there does not exist a matching $\mu^{\prime}\neq\mu$ such that for all players $i\in M\cup W$,
$\mu^{\prime}(i)\neq\mu(i)\ \mbox{implies}\ \mu^{\prime}(i)\ {\succ_{i}}(t_{i}(\tau^{n}(\omega,\mu))\ \mu(i).$
That is, an outcome is persistent Pareto efficient if starting from the outcome for every repetition of the matching there does not exist a Pareto improvement. This notion can be extended to allow for flirting but for simplicity we focus on the case without communication.
We are interested in analogues to the welfare theorems. It is easy to see that the analogue to the first welfare theorem fails. If $(\mu,\omega)$ is a self-confirming outcome, then it may not be persistent Pareto efficient. It may not even be Pareto efficient in the sense of absence of Pareto improvements. Stability of $\mu$ means that there is absence of pairwise common belief in blocking. Yet, as Example 3 demonstrates, there can be cases of self-confirming outcomes where each player of a pair likes to block but this is not pairwise common belief among them, causing the outcome to be Pareto inefficient in the sense that there is a Pareto improvement at $(\mu,\omega)$. Yet, we do have the analogue to the second welfare theorem:
Proposition 8
If $(\mu,\omega)$ is persistent Pareto efficient, then $(\mu,\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu))$ is a self-confirming outcome.
Proof. Any two-sided matching game with unawareness involves a finite number of finite state spaces. Any finite state machine on a finite space must have absorbing sets. Since awareness can only go up, we must have that $\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu)$ is an absorbing state of $\tau$ given $\mu$. Since $(\mu,\omega)$ is persistent Pareto efficient, for any $n$ there is no Pareto improvement $\mu^{\prime}$. Thus, there is also no Pareto improvement at $\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu)$. Hence, there is no blocking pair and thus no pair that has pairwise common belief in blocking. We conclude that $\mu$ is stable at $\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu)$. Since we already observed that $\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu)$ is an absorbing state given $\mu$, we have that $(\mu,\lim_{n\rightarrow\infty}\tau^{n}(\omega,\mu))$ is a self-confirming outcome. $\Box$
Our notion of persistent Pareto efficiency of $(\mu,\omega)$ takes seriously that the players experience $\mu$. Alternatively (or additionally) one may require that also Pareto improvements are experienced. Given an outcome $(\mu,\omega)$, we say that matching $\mu^{\prime}\neq\mu$ is a $\mu^{\prime}$-experienced Pareto improvement if for all players $i\in M\cup W$, $\mu^{\prime}(i)\neq\mu(i)$ implies $\mu^{\prime}(i)\ {\succ_{i}}(t_{i}(\tau(\omega,\mu^{\prime}))\ \mu(i)$. Note that given outcome $(\mu,\omega)$, matching $\mu^{\prime}\neq\mu$ being a $\mu$-experienced Pareto improvement does not imply that $\mu^{\prime}$ is a $\mu^{\prime}$-experienced Pareto improvement and vice versa. For instance, a matching $\mu^{\prime}$ may not be a $\mu$-experienced Pareto improvement but when exposed to $\mu^{\prime}$, players may discover that they like it more than $\mu$ due to unanticipated experiences, awareness changes, and induced preference changes. We believe that a notion of Pareto efficiency that requires absence of $\mu^{\prime}$-experienced Pareto improvements is too demanding and conceptually problematic especially in larger matching markets. It would require experimenting with all matchings.

### v2

@@ Abstract
We study dynamic decentralized two-sided matching where players’ preferences evolve due to unanticipated experiences. Stability requires no pairwise common belief in blocking, but unanticipated experiences can destabilize matchings. We show the existence of self-confirming outcomes that are stable and do not lead to unanticipated experiences. We propose a decentralized matching process that prioritizes mutual optimal blocking pairs with probability $1-\varepsilon$ and picks any other optimal blocking pair with $\varepsilon$, representing market friction. Frictions are necessary for convergence to self-confirming stable outcomes even without unawareness. We extend our results by allowing communication and show convergence to flirt-proof self-confirming outcomes.
@@ 1 Introduction
Many matching problems such as whom to marry, which school to choose, which profession to enter, where to take up a residency etc. involve transformative experiences (Paul, 2014) that change who we are, our beliefs, the things and issues we care about, and our preferences. These preference changes have implications for the stability of matchings. For instance, in the US marriage market, 43% of ever-married couples were divorced or widowed, and 23% of married couples are remarried couples (Livingstone, 2014). During marriage, spouses may become aware of intolerable attributes of the partner that they were previously oblivious to it. Or they may experience events in the family and workplace such as addiction, unemployment, domestic violence, mental health issues etc., some of which can only be fully grasped once experienced. For instance, marriage often involves parenting, itself a transformative experience, with profound changes in preferences as spouses become parents. Svar and Verner (2008) found a negative causal impact of children on relationship duration in Denmark.
@@
In the final Section 7, we discuss the effect of infidelity. We also discuss how players can be confused about other players in self-confirming outcomes leading to (stable) awareness of unawareness (Schipper, 2024). Finally, we discuss the related literature.
@@ 2.3 Random Paths to Stability
Prior section showed that we need to go beyond deterministic processes in order to guarantee stable outcomes in decentralized matching markets. Roth and Vande Vate (1990) showed that we can reach stable outcomes when a process that randomly picks among blocking pairs. In particular, they showed that starting from an arbitrary matching $\mu$, there exist a finite sequence of matchings $\mu_{1},...,\mu_{k}$ such that $\mu=\mu_{1}$ and $\mu_{k}$ is stable, and for $\mu_{i}$, $i=1,...,k-1$, there is a blocking pair $(m_{i},w_{i})$ that if satisfied yields matching $\mu_{i+1}$. As a corollary, they show that the process of satisfying randomly chosen blocking pairs will converge to a stable matching with probability 1.
@@
This observation has a profound economic corollary: Frictions that prevent the satisfaction of mutually optimal blocking pairs are necessary for any decentralized matching processes to reach stability. This observation runs counter to standard economic wisdom that frictions hamper the functioning of decentralized markets. In contrast, for decentralized matching markets, small frictions are necessary for reaching stable matchings. In a separate online-appendix, we present another application of our cyclic example to entry in matching markets.
@@ 2.4 Arbitrarily Small Frictions are Enough
For the unperturbed process, both stable matchings and cycles formed by satisfying unique mutually optimal blocking pairs are absorbing sets. However, we argue that for an arbitrarily small $\varepsilon>0$, which captures frictions that prevent the matching of mutually optimal blocking pairs, the perturbed process converges to a stable matching in finite time with probability 1. Given a marriage market $\langle M,W,(\succ_{i})_{i\in M\cup W}\rangle$, cycle(s) formed by satisfying unique mutually optimal blocking pairs at each step may or may not exist. If there is no such cycle, our claim is trivially true. If there exists such a cycle, then conditional on entering such a cycle, the probability that the process stays in this cycle for exactly $k$ period is $(1-\varepsilon)^{k-1}\varepsilon$. Hence the probability that the process stays in this cycle forever is $\lim_{k\rightarrow\infty}(1-\varepsilon)^{k-1}\varepsilon=0$, which means that the process leaves the cycle in finite time with probability 1. After leaving the cycle, the process might come back to this cycle or enter another cycle. However, the same argument applies, and the process leaves the cycle again in finite time with probability 1. As soon as the process catches one of Ackermann et al. (2008)’s optimal blocking pairs, it is trapped in a basin of attraction of a stable matching. Note that by the $\varepsilon$-events, any optimal blocking pair has a strict positive probability. Thus, we can make use of the result by Ackermann et al. (2008) according to which for any unstable matching, there exist a finite sequence of satisfying optimal blocking pairs that leads to a stable matching. The process may potentially spend long periods in the cycle(s), but only the stable matchings are absorbing
@@ 3 Matching under Unawareness
We illustrate an unawareness structure with point-beliefs for two players satisfying Assumptions 1 and 2 in Figure 3. There are two characteristics of players or experiences with players of which players may or may not be unaware. Thus, we have four spaces. The left space $S_{1}$ models the situation in which only the first characteristic/experience is expressible, while the right space $S_{2}$ models the situations when only the second characteristic/experience is expressible. On the upmost space $S_{3}$ both characteristics/experiences are expressible. In the lowest space $S_{0}$ none of the characteristics/experiences are expressible. States model combinations of awareness of both agents in a consistent way as described by the type mapping and the conditions imposed on the type mapping. The type mapping of one player (i.e., the man) is is indicated with blue solid lines while the one of the other player is indicated with pink dash lines (i.e., the woman). For instance, at state $\omega_{10}$ the woman is aware of both characteristics/experiences, because her type mapping maps $\omega_{10}$ to itself, while the man is only aware of the first characteristic/experience, because his type mapping maps $\omega_{10}$ to $\omega_{4}$ in $S_{1}$. At $\omega_{4}$ both players are aware of the first characteristic/experience. Thus, at $\omega_{10}$ the man point-believes that the woman is aware of the first characteristics/experiences, which is consistent with his awareness at $\omega_{10}$ since he can only envision the first characteristic/experience but not the second at $\omega_{10}$. More interesting is for instance state $\omega_{24}$. At that state, the man is aware of the first characteristic/experience only while the woman is only aware of the second characteristics/experiences only because their type mappings map to $\omega_{3}$ and $\omega_{6}$, respectively. At $\omega_{3}$, the man correctly believes that the woman is unaware of characteristic/experience 1 while at $\omega_{6}$ the woman correctly believes that the man is unaware of characteristic/experience 2. This is indicated by the type mappings that map the respective states to $\omega_{0}$. We omit the projections in order not to clutter the figure further, but they should be clear. For instance, states $\omega_{13},\omega_{14}$ and $\omega_{21}$ all project to $\omega_{7}$. Assumptions 1 and 2 are easy to verify for instance in spaces $S_{1}$ and $S_{2}$ (albeit less obvious in $S_{3}$). Note that any unawareness structure for correct point-beliefs of two players and two characteristics/experiences satisfying Assumptions 1 and 2 must look like Figure 3. With more players, which we need in non-trivial matching games, or more characteristics/experiences, they become more complex. However, in our examples we will often just focus on the states that are relevant for the feature we aim to illustrate with the example.
@@ 4 Self-Confirming Stable Outcomes
Stability itself is not a satisfactory solution in our setting because $\mu$ being stable at $\omega$ does not rule out that some players discover something in the matching $\mu$ at $\omega$ that changes their preferences and consequently destabilizes the previously stable matching so that they want to get divorced. For a satisfactory solution, we also need that beliefs are stable given the matching.
@@
We define the matching process by transition probabilities on the product space $\Omega\times\mathsf{M}$. This warrants some explanations. Naturally, the process would proceed as follows: Given a current state $\omega$ and awareness/preferences at $\omega$, the current match $\mu$ may feature some blocking pairs w.r.t. preferences at $\omega$. Satisfying some blocking pair (i.e., preferably some mutual optimal blocking pair) would lead to another match $\mu^{\prime}$. At this match $\mu^{\prime}$ and state $\omega$, there may be discoveries leading to another state $\tau(\omega,\mu^{\prime})=\omega^{\prime}$. At this state and corresponding awareness/preferences, there might be blocking pairs. Satisfying a blocking pair may lead to yet another matching $\mu^{\prime\prime}$ While in this explanation, we let states transit after the transition of the matches, the above formal description of the process involves a simultaneous transition of states and matches. It can be thought of just describing the process at every “even” period. This is well-defined because states move deterministically according to a finite state machine given the prior state and the new match. The process on $\Omega\times\mathsf{M}$ is not deterministic though because blocking pairs are selected randomly with $(1-\varepsilon)$ assigned to mutual best blocking pairs if they exist.
@@ 5 Allowing for Flirting
$H_{m}(\omega,\mu,w):=\{\omega^{\prime}\in S_{t_{m}(\omega)}:\omega^{\prime}\in CK_{m,w}([m\succ_{w}\mu(w)]\cap[w\succ_{m}\mu(m)]),S_{t_{w}(\omega^{\prime})}\trianglerighteq S_{t_{w}(t_{m}(\omega))}\}$
@@ 6 Does Divorce Improve Welfare?
Example 8 Suppose there are five men $M=\{m_{1},m_{2},m_{3},m_{4},m_{5}\}$ and five women $W=\{w_{1},w_{2},w_{3},w_{4},w_{5}\}$. Consider the following strict preferences given by the rank order lists:
@@
and the initial stable matching $\mu_{1}=\begin{pmatrix}m_{1}&m_{2}&m_{3}&m_{4}&m_{5}\\ w_{1}&w_{2}&w_{3}&w_{4}&w_{5}\end{pmatrix}$.
@@
Suppose $m_{1}$ is enlightened and changes his preference such that his new rank order list is $w_{2},w_{1},w_{4},w_{3},w_{5}$. For simplicity, suppose that this is the only change in the players’ preferences along the entire process and that all players have correct beliefs about every other agents’ preferences. With $m_{1}$’s new preference, he blocks with $w_{2}$. We show that if $m_{1}$ rematches to $w_{2}$, the process of satisfying mutually optimal blocking pairs may lead to any of the three stable matchings, depending on which mutually optimal blocking pair is satisfied at stages 7 or 8 (see Figure 12). Furthermore, in these stable matchings, $m_{1}$ can be matched to $w_{2}$ (in $\mu_{10_{A}}$), $w_{1}$ (in $\mu_{11_{B}}$), or $w_{4}$ (in $\mu_{9}$), which means that he may be better off, same, or worse off.